OpenStudy (anonymous):

a driver shifts into neutral when her 1500 kg car is moving at 30 m/s and finds that her speed has dropped 25 m/s ten seconds later. How great was the force of friction which acted to slow her car?

OpenStudy (abhisar):

Hello @dray71837 , Welcome to openstudy !

OpenStudy (abhisar):

Do u have the answer to this question ?

OpenStudy (abhisar):

Initial speed = 30 m/s Final speed = 30-25=5 m/s Time = 10s Using v=u-at 5=30-10a a=2.5 m/s^2 Force= mass*acceleration Force=1500*2.5=37500N

OpenStudy (abhisar):

*3750N

OpenStudy (anonymous):

Thank you very much!